In consider the function 1/x

Find the slope of the line L tangent to the graph of (3,f(3))
I know to find the slope of the tangent I use: lim h->0 f(x+h)-f(x)/h
why does this switch from that to (1/(x+h))-(1/x)/(h)?
I thought since we have x and y values (3,f(3)) I can just somehow plug those values in. Why does x+h become 1/(x+h)
Do we first have to alter the limit definition of a derivative (tangent slope) first given the graph shape? then do some rearranging and factoring? Because I know it gets to be lim h->0 1/x^2 which makes the slope m = -1/9
what I also don't get is why do y become y=1/3 when the original points were (3,f(3))
Confused so much about all this.
Please help
Thank you

Answer by rothauserc(4718)

We are working with the limit definition of the first derivative, that is
Limit as h --》0 of ( f(x+h) - f(x) ) / h
We are given f(x) = 1 / x
We apply the given function to the limit definition
Limit as h --》0 of ( 1/(x+h) - 1/x ) / h =
( x - x - h ) / (x(x+h)h ) =
-h / (x^2+xh)h =
-1 / x^2
Note that the xh term becomes 0 as h --》0
f(3) = 1 / 3
m = f'(3) = -1 / 3^2 = -1 / 9